# Five-Term Relations for wreath Macdonald polynomials and tableau formulas for Pieri coefficients

### v1

@@ 1. Introduction
Tesler’s identity and the five-term relation of Garsia and Mellit provide two of the most powerful tools in developing symmetric function identities related to modified Macdonald polynomials. From here, one can give a number of consequences, such as proving tableau formulas for Pieri coefficients [BH13] or giving identities for Theta operators [DR23], to name a few relevant to this work. The recursion for Pieri coefficients implied by the five-term relation gives a fast, if not the fastest, way of computing the monomial expansion for modified Macdonald polynomials.
A version of Tesler’s identity was recently proved for wreath Macdonald polynomials in [RW25], where a number of plethystic operators were also developed. Here, we revisit the plethystic calculus on multisymmetric functions, giving a concise way of notating vector (or matrix) plethysms. We will demonstrate how to utilize this device by reproving reciprocity identities from the wreath Tesler identitity.
We will then give several families of five-term relations for operators on wreath Macdonald polynomials, and as a consequence, give formulas for computing Pieri coefficients in terms of tableaux. We will also introduce a wreath version for the Theta operators of D’Adderio, Iraci, and Vanden Wyngaerd [DIVW21], and show how they relate to the $\mathbb{D}$ operators in [RW25].
@@ 4. Wreath Macdonald polynomials and the Cauchy kernel
$\langle F,G\rangle_{\ast}=\langle F[\iota X^{\bullet}],G[M^{T}X^{\bullet}]\rangle.$
@@ 4.1. An extension of multisymmetric functions
We let $\mathcal{W}=\Lambda^{\otimes r}\otimes\mathbb{C}[Q]$, where $\mathbb{C}[Q]$ is the span of elements $e^{(c_{0},\dots,c_{r-1})}$ with $c_{0}+\cdots+c_{r-1}=0$ (that is, an element of the weight lattice for $A_{r-1}$). Such an element $(c_{0},\dots,c_{r-1})$ corresponds to an $r$-core, say $\alpha$, for which we instead write $e^{(c_{0},\dots,c_{r-1})}=e^{\alpha}$.
@@
$\left\langle F\otimes e^{\alpha},G\otimes e^{\beta}\right\rangle_{\ast}\coloneqq\langle F,G\rangle_{\ast}\delta_{w_{0}\alpha,\beta}$ and $\left\langle F\otimes e^{\alpha},G\otimes e^{\beta}\right\rangle\coloneqq\langle F,G\rangle\delta_{\alpha,\beta}.$
@@ 5. Operators on wreath Macdonald polynomials
$G\left[\epsilon_{i}\frac{D}{M}\right]H_{\lambda}\otimes e^{\mathrm{core}(\alpha)}=G\left[\epsilon_{i}\left(\frac{D_{\lambda}}{M_{0}}\right)^{(i)}\right]H_{\lambda}\otimes e^{\mathrm{core}(\alpha)}.$
@@ 6. The Tesler Identity and the map $\mathsf{V}$
$\operatorname{\mathcal{T}}^{(i)}_{v}\mathsf{V}=\mathsf{V}\Omega\left[-v\epsilon_{i}\frac{D}{M}\right]$ and $\mathsf{V}\mathcal{P}_{-v/M^{T}}^{(i)}=\Omega\left[-v\epsilon_{i}\frac{\iota D^{\dagger}}{M}\right]\mathsf{V}.$
@@ 6.1. Wreath Macdonald-Koornwinder reciprocity revisited
    $=\nabla^{-1}\Omega\left[-z\epsilon_{0}\frac{D^{\dagger}}{M}\right]H_{\lambda}[\iota(\epsilon_{0}+z\iota D^{\dagger})]\mathsf{V}\left(1\right)$
@@
$\frac{H_{\lambda}[\epsilon_{0}+z\iota D_{\mu}^{\bullet}]}{\prod_{\begin{subarray}{c}\square\in\mu\\ \bar{c}_{\square}=0\end{subarray}}1-z\chi_{\square}}=\frac{H_{\mu}[\epsilon_{0}+z\iota D_{\lambda}^{\bullet}]}{\prod_{\begin{subarray}{c}\square\in\lambda\\ \bar{c}_{\square}=0\end{subarray}}1-z\chi_{\square}}.$
@@
$\frac{H_{\lambda}[\epsilon_{0}+z\sigma^{k}\iota D_{\mu}^{\bullet}]}{\prod_{\begin{subarray}{c}\square\in\mu\\ \bar{c}_{\square}=k\end{subarray}}1-z\chi_{\square}}=\frac{H_{\mu}[\epsilon_{0}+z\sigma^{k}\iota D_{\lambda}^{\bullet}]}{\prod_{\begin{subarray}{c}\square\in\lambda\\ \bar{c}_{\square}=k\end{subarray}}1-z\chi_{\square}}.$
@@
$\frac{H_{\lambda}[\epsilon_{0}-\epsilon_{-k}+\sigma^{k}\iota M^{T}B_{\mu}^{\bullet}]}{\Pi_{\mu}^{(k)}}=\frac{H_{\mu}[\epsilon_{0}-\epsilon_{-k}+\sigma^{k}\iota M^{T}B_{\lambda}^{\bullet}]}{\Pi_{\lambda}^{(k)}}.$
@@
$\frac{H_{\lambda}[M\iota B_{\mu}^{\bullet}]}{\Pi_{\mu}^{(0)}}=\frac{H_{\mu}[M\iota B_{\lambda}^{\bullet}]}{\Pi_{\lambda}^{(0)}}.$
@@ 6.2. For empty cores
$H_{\lambda}\left[\epsilon_{0}-z\epsilon_{-k}\right]=\prod_{\begin{subarray}{c}\square\in\lambda\\ \bar{c}_{\square}=k\end{subarray}}1-z\chi_{\square}.$
@@
$\langle H_{\lambda},h_{n-j}[X^{(0)}]e_{j}[X^{(-k)}]\rangle=e_{j}\left[B^{(k)}_{\lambda}\right].$
@@
$\left\langle e_{j}[\epsilon_{k}B]~{}F\otimes e^{\alpha},h_{n}[X^{(0)}]\otimes e^{\alpha}\right\rangle=\left\langle F,h_{n-j}[X^{(0)}]e_{j}[X^{(-k)}]\right\rangle.$
@@
$\prod_{\begin{subarray}{c}\square\in\lambda\\ \bar{c}_{\square}=k\end{subarray}}1-z\chi_{\square}$ $=H_{\lambda}[\epsilon_{0}-z\epsilon_{-k}]$
$=\sum_{i+j=n}h_{i}[X^{(0)}]^{\perp}(-z)^{j}e_{j}[X^{(-k)}]^{\perp}H_{\lambda}$
$=\left\langle H_{\lambda},\sum_{j=0}^{n}h_{n-j}[X^{(0)}](-z)^{j}e_{j}[X^{(-k)}]\right\rangle.$
@@
It is generally unclear at this point how to compute $H_{\lambda}[-\epsilon_{0}]$ for general $\lambda$. However, one of the consequences of the formulas for Pieri coefficients is giving a tableau formula for this evaluation. In [OS23], the authors present an unpublished result of Haiman that describes $H_{\lambda}[-\epsilon_{i}]$.
@@ 7.1. Triangularities in $u$ and $v$
$\mathsf{V}^{-1}\mathsf{V}\mathcal{P}^{(p)}_{u/M^{T}}\Delta_{v}^{(s)}\mathcal{P}^{(p)}_{-u/M^{T}}$ $=\mathsf{V}^{-1}\Omega\left[u\epsilon_{p}\frac{\iota D^{\dagger}}{M}\right]\operatorname{\mathcal{T}}_{v}^{(s)}\Omega\left[-u\epsilon_{p}\frac{\iota D^{\dagger}}{M}\right]\mathsf{V}.$
@@
$\Omega\left[u\epsilon_{p}\frac{\iota D}{M}\right]~{}h_{k}^{\perp}\left[X^{(s)}\right]\Omega\left[-u\epsilon_{p}\frac{\iota D}{M}\right]$
is a polynomial in $u$ of degree $k$. Therefore, the degree of $u$ in $\mathcal{P}^{(p)}_{u/M}\Delta_{v}^{(s)}\mathcal{P}^{(p)}_{-u/M}$ is at most the degree of $v$. Since the operator $\Omega[v\epsilon_{s}(D/M)]$ only increases the degree in $u$, we have that $W_{i,j}=0$ for $j<i$. ∎
@@ 7.3. The dual Pieri formulas
$\downarrow\nabla=\nabla^{-1}\downarrow,$ $\downarrow\operatorname{\mathcal{T}}^{(i)}_{A}=\operatorname{\mathcal{T}}^{(-i)}_{-A^{T}_{\ast}}\downarrow,$ $\downarrow\mathcal{P}^{(i)}_{A}=\mathcal{P}^{(-i)}_{-A^{T}_{\ast}}\downarrow,$ $\downarrow F[\epsilon_{s}D^{\dagger}]=F[\epsilon_{s}D_{\ast}]\downarrow.$
@@
$\mathsf{V}_{\ast}=\nabla^{-1}\mathcal{P}_{1/M^{T}_{\ast}}^{(0)}\operatorname{\mathcal{T}}_{-1}^{(0)}\nabla^{-1}.$
@@
${}^{\ast}\Delta_{v}^{(i)}\coloneqq\Omega\left[-v\epsilon_{i}\frac{qtD_{\ast}}{M}\right]$ and its inverse ${}^{\ast}\Delta_{-v}^{(i)}\coloneqq\Omega\left[v\epsilon_{i}\frac{qtD_{\ast}}{M}\right].$
@@
${}^{\ast}\Delta_{-v}^{(s)}h_{k}^{\perp}[X^{(p)}]{}^{\ast}\Delta_{v}^{(s)}H_{\mu}\otimes e^{\alpha}$ $=\sum_{\lambda\subset_{k}\mu}c_{\lambda,\mu}^{(p)}H_{\lambda}\Omega\left[v\epsilon_{s}\frac{qt(D_{\lambda}-D_{\mu})_{\ast}}{M}\right]$
@@
$\operatorname{\mathcal{T}}_{-u}^{(p)}{}^{\ast}\Delta_{-v}^{(s)}\operatorname{\mathcal{T}}^{(p)}_{u}\mathsf{V}_{\ast}=\mathsf{V}_{\ast}\Omega\left[-u\epsilon_{p}\frac{qt\iota D^{\dagger}_{\ast}}{M}\right]\mathcal{P}_{vqtX/M}^{(s)}\Omega\left[u\epsilon_{p}\frac{qt\iota D^{\dagger}_{\ast}}{M}\right].$
@@
${}^{\ast}\Delta_{-v/u}^{(s)}\operatorname{\mathcal{T}}_{u}^{(p)}{}^{\ast}\Delta_{v/u}^{(s)}H_{\mu}\otimes e^{\alpha}\Big{|}_{u^{l}v^{k}}=\sum_{\lambda\subset_{k+l}\mu}c_{\mu,\lambda}^{(p)}\Omega\left[-v\epsilon_{i}\frac{qt(D_{\mu}-D_{\lambda})_{\ast}}{M}\right]\Big{|}_{v^{k}}H_{\lambda}\otimes e^{\alpha}.$
@@
$c_{\mu,\lambda}^{(p)}=\frac{(-1)^{k}}{e_{k}\left[\left(B^{(s)}_{\mu/\lambda}\right)_{\ast}\right]}\sum_{\lambda\subset_{l}\nu\subset_{k}\mu}\frac{1}{T_{\mu/\nu}^{(s)}}c_{\mu,\nu}^{(p)}c_{\nu,\lambda}^{(p)}.$
@@
Therefore, Corollary 7.10 gives a tableau formula for the monomial expansion of wreath Macdonald polynomials.

### v3

@@ 1. Introduction
Defined by Haiman [Hai03], the wreath Macdonald polynomials generalize the modified Macdonald polynomials from the symmetric group $\Sigma_{n}$ to its wreath product $\Gamma_{n}:=\mathbb{Z}/r\mathbb{Z}\wr\Sigma_{n}$. Just as one can study the representation theory of the symmetric groups $\{\Sigma_{n}\}$ using the ring of symmetric functions $\Lambda$, for fixed $r$, one can model the representation theory of $\{\Gamma_{n}\}$ via the ring of multisymmetric functions $\Lambda^{\otimes r}$ [Mac15]. The wreath Macdonald polynomials are distinguished elements of the deformed ring $\mathbb{C}(q,t)\otimes\Lambda^{\otimes r}$. However, it may be naive to simply treat wreath Macdonald theory as a partially symmetric analogue of Macdonald theory. Haiman’s original definition was motivated by geometry. Just as his celebrated proof of the Macdonald positivity conjecture [Hai01] showed that Macdonald polynomials are intimately tied to the geometry of Hilbert schemes of points on $\mathbb{C}^{2}$, Haiman conjectured that the wreath Macdonald polynomials should be related to certain Nakajima quiver varieties for the cyclic quiver. This conjecture was realized by Bezrukavnikov–Finkelberg [BF14], who proved a wreath analogue of Macdonald positivity.
Functional aspects of wreath Macdonald polynomials did not receive much attention until recently, and it was not evident which properties and identities had wreath analogues. For instance, the usual Macdonald polynomials are solutions to $q$-difference equations; wreath analogues of this were finally established by the second author with Orr and Shimozono [Wen25a, OSW22]. More relevant to this paper, Tesler’s identity and the five-term relation of Garsia and Mellit [GM19] provide two of the most powerful tools in developing symmetric function identities related to modified Macdonald polynomials, even finding applications in geometry. From here, one can give a number of consequences, such as proving tableau formulas for Pieri coefficients [BH13] or giving identities for Theta operators [DR23]. The recursion for Pieri coefficients implied by the five-term relation gives a fast, if not the fastest, way of computing the monomial expansion for modified Macdonald polynomials. A version of Tesler’s identity was recently proved for wreath Macdonald polynomials in [RW25], where a number of plethystic operators were also developed.
In this paper, we derive consequences of the wreath Tesler identity and push the functional theory of wreath Macdonald polynomials further into maturation. First, we revisit the plethystic calculus on multisymmetric functions, giving a concise way of notating vector (or matrix) plethysms. We then demonstrate how to utilize this device by reproving reciprocity identities from the wreath Tesler identity. Next, we give several families of five-term relations for operators on wreath Macdonald polynomials, and as a consequence, give formulas for computing Pieri coefficients in terms of tableaux. To clarify, our Pieri rules are not the wreath analogues of those found in [Mac15, Chapter VI]; it appears that such a wreath analogue is quite complicated (cf. [Wen25b]). Nonetheless, these recursive formulas are incredibly powerful, allowing one to compute many more examples of wreath Macdonald polynomials than was previously possible.
We also introduce a wreath version for the Theta operators of D’Adderio, Iraci, and Vanden Wyngaerd [DIVW21], and show how they relate to the $\mathbb{D}$ operators in [RW25]. In the case of modified Macdonald polynomials, Theta operators provide a framework for producing many useful identities [DR23]. They were originally introduced to give a compositional refinement of the Delta Conjecture of Haglund, Remmel, and Wilson [HRW18]. This refinement was proved by D’Adderio and Mellit [DM22]. Theta operators also conjecturally produce the Frobenius characteristic for the space of coinvariants in two sets of commuting and two sets of anticommuting variables [DIVW21]. We predict that wreath Theta operators will also give a powerful tool for dealing with multisymmetric function identities related to wreath Macdonald polynomials. Here, we will only look at an identity that explains the relation between wreath Theta operators and the action of the quantum toroidal algebra on multisymmetric functions.
@@ 4. Wreath Macdonald polynomials and the Cauchy kernel
$\langle F,G\rangle_{\ast}=\langle F[-\iota X^{\bullet}],G[M^{T}X^{\bullet}]\rangle.$
@@ 4.1. An extension of multisymmetric functions
We let $\mathcal{W}=\Lambda^{\otimes r}\otimes\mathbb{C}[Q]$, where $\mathbb{C}[Q]$ is the span of elements $e^{(c_{0},\dots,c_{r-1})}$ with $c_{0}+\cdots+c_{r-1}=0$ (that is, an element of the root lattice for $A_{r-1}$). Such an element $(c_{0},\dots,c_{r-1})$ corresponds to an $r$-core, say $\alpha$, for which we instead write $e^{(c_{0},\dots,c_{r-1})}=e^{\alpha}$.
@@
$\left\langle F\otimes e^{\alpha},G\otimes e^{\beta}\right\rangle_{\ast}\coloneqq\langle F,G\rangle_{\ast}\delta_{w_{0}\alpha,\beta}$ and $\left\langle F\otimes e^{\alpha},G\otimes e^{\beta}\right\rangle\coloneqq\langle F,G\rangle\delta_{w_{0}\alpha,\beta}.$
@@ 5. Operators on wreath Macdonald polynomials
$=\epsilon_{i}\left(\sum_{p\in\mathbb{Z}/r\mathbb{Z}}\epsilon_{p}\sum_{a,b\geq 0}q^{a}t^{b}D^{(p+a-b)}_{\lambda}\right)$
@@
$G\left[\epsilon_{i}\frac{D}{M}\right]H_{\lambda}\otimes e^{\mathrm{core}(\alpha)}=G\left[\epsilon_{i}\left(\frac{D_{\lambda}}{M_{0}}\right)^{(i)}\right]H_{\lambda}\otimes e^{\mathrm{core}(\alpha)}.$ (5.1)
@@ 6. The Tesler Identity and the map $\mathsf{V}$
$\operatorname{\mathcal{T}}^{(i)}_{v}\mathsf{V}=\mathsf{V}\Omega\left[-v\epsilon_{i}\frac{D}{M}\right]$ and $\mathsf{V}\mathcal{P}_{-v/M^{T}}^{(i)}=\Omega\left[-v\epsilon_{i}\frac{\iota D^{\dagger}}{M^{T}}\right]\mathsf{V}.$
@@ 6.1. Wreath Macdonald-Koornwinder reciprocity revisited
    $=\nabla^{-1}\Omega\left[-z\epsilon_{0}\frac{\iota D^{\dagger}}{M^{T}}\right]H_{\lambda}[\iota(\epsilon_{0}+z\iota D^{\dagger})]\mathsf{V}\left(1\right)$
@@
$\frac{H_{\lambda}[\epsilon_{0}+z\iota D_{\mu}^{\bullet}]}{\prod_{\begin{subarray}{c}\square\in\lambda\\ \bar{c}_{\square}=0\end{subarray}}1-z\chi_{\square}}=\frac{H_{\mu}[\epsilon_{0}+z\iota D_{\lambda}^{\bullet}]}{\prod_{\begin{subarray}{c}\square\in\mu\\ \bar{c}_{\square}=0\end{subarray}}1-z\chi_{\square}}.$
@@
$\frac{H_{\lambda}[\epsilon_{0}+z\iota\sigma^{k}D_{\mu}^{\bullet}]}{\prod_{\begin{subarray}{c}\square\in\lambda\\ \bar{c}_{\square}=k\end{subarray}}1-z\chi_{\square}}=\frac{H_{\mu}[\epsilon_{0}+z\iota\sigma^{k}D_{\lambda}^{\bullet}]}{\prod_{\begin{subarray}{c}\square\in\mu\\ \bar{c}_{\square}=k\end{subarray}}1-z\chi_{\square}}.$ (6.1)
Remark 6.4.
The plethysms in (6.1) look different from the ones from [RW25]. This is due to the different treatment of plethysm in this paper. The two formalisms coincide in their action on multi-symmetric functions. On the other hand, in loc. cit., we do not define $\sigma$ and $\iota$ on a sum of characters like $D_{\mu}^{\bullet}$; rather, an expression such as $f[\sigma^{k}\iota D_{\mu}^{\bullet}]$ is meant to be interpreted as performing $f[\sigma^{k}\iota X^{\bullet}]$ and then evaluating at $D_{\mu}^{\bullet}$. This results in the sequence of transformations
$p_{n}[X^{(i)}]\mapsto p_{n}[X^{(k-i)}]\mapsto p_{n}[D_{\mu}^{(k-i)}].$
In the formalism used in this paper, we have
$\epsilon_{i}\left(\iota\sigma^{k}D_{\mu}^{\bullet}\right)=D_{\mu}^{(k-i)}.$
Therefore, the evaluations in [RW25] and (6.1) coincide.
@@
$\frac{H_{\lambda}[\epsilon_{0}-\epsilon_{k}+\iota\sigma^{k}M^{T}B_{\mu}^{\bullet}]}{\Pi_{\lambda}^{(k)}}=\frac{H_{\mu}[\epsilon_{0}-\epsilon_{k}+\iota\sigma^{k}M^{T}B_{\lambda}^{\bullet}]}{\Pi_{\mu}^{(k)}}.$
@@
$\frac{H_{\lambda}[M\iota B_{\mu}^{\bullet}]}{\Pi_{\lambda}^{(0)}}=\frac{H_{\mu}[M\iota B_{\lambda}^{\bullet}]}{\Pi_{\mu}^{(0)}}.$
@@ 6.2. For empty cores
$H_{\lambda}\left[\epsilon_{0}-z\epsilon_{k}\right]=\prod_{\begin{subarray}{c}\square\in\lambda\\ \bar{c}_{\square}=k\end{subarray}}1-z\chi_{\square}.$
@@
$\langle H_{\lambda},h_{n-j}[X^{(0)}]e_{j}[X^{(k)}]\rangle=e_{j}\left[B^{(k)}_{\lambda}\right].$
@@
$\left\langle e_{j}[\epsilon_{k}B]~{}F\otimes e^{\alpha},h_{n}[X^{(0)}]\otimes e^{\alpha}\right\rangle=\left\langle F,h_{n-j}[X^{(0)}]e_{j}[X^{(k)}]\right\rangle.$
@@
$\prod_{\begin{subarray}{c}\square\in\lambda\\ \bar{c}_{\square}=k\end{subarray}}1-z\chi_{\square}$ $=H_{\lambda}[\epsilon_{0}-z\epsilon_{k}]$
$=\sum_{i+j=n}h_{i}[X^{(0)}]^{\perp}(-z)^{j}e_{j}[X^{(k)}]^{\perp}H_{\lambda}$
$=\left\langle H_{\lambda},\sum_{j=0}^{n}h_{n-j}[X^{(0)}](-z)^{j}e_{j}[X^{(k)}]\right\rangle.$
@@
It is generally unclear at this point how to compute $H_{\lambda}[-\epsilon_{0}]$ for general $\lambda$. However, one of the consequences of the formulas for Pieri coefficients is giving a tableau formula for this evaluation. In [OS24], the authors present an unpublished result of Haiman that describes $H_{\lambda}[-\epsilon_{i}]$.
@@ 7. Five-term relations and tableau formulas for Pieri coefficients
We see from (5.1) that
$\Delta_{v}^{(s)}H_{\lambda}=\Omega\left[-v\left(\frac{D_{\lambda}}{M_{0}}\right)^{(s)}\right]H_{\lambda}.$
@@ 7.1. Triangularities in $u$ and $v$
$\mathsf{V}^{-1}\mathsf{V}\mathcal{P}^{(p)}_{u/M^{T}}\Delta_{v}^{(s)}\mathcal{P}^{(p)}_{-u/M^{T}}$ $=\mathsf{V}^{-1}\Omega\left[u\epsilon_{p}\frac{\iota D^{\dagger}}{M^{T}}\right]\operatorname{\mathcal{T}}_{v}^{(s)}\Omega\left[-u\epsilon_{p}\frac{\iota D^{\dagger}}{M}\right]\mathsf{V}.$
@@
$\Omega\left[u\epsilon_{p}\frac{\iota D^{\dagger}}{M^{T}}\right]~{}h_{k}^{\perp}\left[X^{(s)}\right]\Omega\left[-u\epsilon_{p}\frac{\iota D^{\dagger}}{M^{T}}\right]$
is a polynomial in $u$ of degree $k$. Therefore, the degree of $u$ in $\mathcal{P}^{(p)}_{u/M^{T}}\Delta_{v}^{(s)}\mathcal{P}^{(p)}_{-u/M^{T}}$ is at most the degree of $v$. Since the operator $\Omega[v\epsilon_{s}(D/M)]$ only increases the degree in $u$, we have that $W_{i,j}=0$ for $j<i$. ∎
@@ 7.3. The dual Pieri formulas
$\downarrow\nabla\downarrow=\nabla^{-1},$ $\downarrow\operatorname{\mathcal{T}}^{(i)}_{A}\downarrow=\operatorname{\mathcal{T}}^{(i)}_{-\iota A_{\ast}},$ $\downarrow\mathcal{P}^{(i)}_{A}\downarrow=\mathcal{P}^{(-i)}_{-A_{\ast}\iota},$ $\downarrow F[\epsilon_{s}D^{\dagger}]\downarrow=F[\epsilon_{s}D_{\ast}].$
@@
$\mathsf{V}_{\ast}=\nabla^{-1}\mathcal{P}_{1/M_{\ast}}^{(0)}\operatorname{\mathcal{T}}_{-1}^{(0)}\nabla^{-1}.$
@@
${}^{\ast}\Delta_{v}^{(i)}\coloneqq\Omega\left[-v\epsilon_{i}\frac{D_{\ast}}{M_{\ast}}\right]$ and its inverse ${}^{\ast}\Delta_{-v}^{(i)}\coloneqq\Omega\left[v\epsilon_{i}\frac{D_{\ast}}{M_{\ast}}\right].$
In particular, note that since
$\epsilon_{i}\frac{(D_{\lambda}^{\bullet})_{\ast}}{M_{\ast}}$ $=\epsilon_{i}\sum_{a,b\geq 0}q^{-a}t^{-b}\sigma^{a-b}\sum_{j\in\mathbb{Z}/r\mathbb{Z}}\epsilon_{j}(D^{(j)}_{\lambda})_{\ast}$
$=\epsilon_{i}\sum_{a,b\geq 0}q^{-a}t^{-b}\sum_{j\in\mathbb{Z}/r\mathbb{Z}}\epsilon_{j-a+b}(D^{(j)}_{\lambda})_{\ast}$
$=\epsilon_{i}\sum_{p\in\mathbb{Z}/r\mathbb{Z}}\epsilon_{p}\left(\sum_{a,b\geq 0}q^{a}t^{b}D^{(p+a-b)}_{\lambda}\right)_{\ast}$
$=\epsilon_{i}\left(\frac{D_{\lambda}}{M_{0}}\right)^{(i)}_{\ast},$
we have
${}^{\ast}\Delta_{v}^{(i)}H_{\lambda}=\Omega\left[-v\left(\frac{D_{\lambda}}{M_{0}}\right)^{(i)}_{\ast}\right]H_{\lambda}.$
@@
${}^{\ast}\Delta_{-v}^{(s)}h_{k}^{\perp}[X^{(p)}]{}^{\ast}\Delta_{v}^{(s)}H_{\mu}\otimes e^{\alpha}$ $=\sum_{\lambda\subset_{k}\mu}c_{\lambda,\mu}^{(p)}H_{\lambda}\Omega\left[v\epsilon_{s}\frac{(D_{\lambda}^{\bullet}-D_{\mu}^{\bullet})_{\ast}}{M_{\ast}}\right]$
@@
$\operatorname{\mathcal{T}}_{-u}^{(p)}{}^{\ast}\Delta_{-v}^{(s)}\operatorname{\mathcal{T}}^{(p)}_{u}\mathsf{V}_{\ast}=\mathsf{V}_{\ast}\Omega\left[-u\epsilon_{p}\frac{qt\iota D^{\dagger}_{\ast}}{M}\right]\mathcal{P}_{v/M_{\ast}}^{(s)}\Omega\left[u\epsilon_{p}\frac{qt\iota D^{\dagger}_{\ast}}{M}\right].$
@@
${}^{\ast}\Delta_{-v/u}^{(s)}\operatorname{\mathcal{T}}_{u}^{(p)}{}^{\ast}\Delta_{v/u}^{(s)}H_{\mu}\otimes e^{\alpha}\Big{|}_{u^{l}v^{k}}=\sum_{\lambda\subset_{k+l}\mu}c_{\mu,\lambda}^{(p)}\Omega\left[-v\epsilon_{i}\left(\frac{D_{\mu}-D_{\lambda}}{M_{0}}\right)^{(i)}_{\ast}\right]\Big{|}_{v^{k}}H_{\lambda}\otimes e^{\alpha}.$
@@
$c_{\lambda,\mu}^{(p)}=\frac{(-1)^{k}}{e_{k}\left[\left(B^{(s)}_{\mu/\lambda}\right)_{\ast}\right]}\sum_{\lambda\subset_{l}\nu\subset_{k}\mu}\frac{1}{T_{\mu/\nu}^{(s)}}c_{\nu,\mu}^{(p)}c_{\lambda,\nu}^{(p)}.$
@@
Therefore, Theorem 7.10 gives a tableau formula for the monomial expansion of wreath Macdonald polynomials.